# size of isogeny classes of abelian varieties in the extended lubin-tate case

### v1

@@ Abstract.
We prove a lower bound for the size of the isogeny class of a simple abelian variety over a finite field with commutative endomorphism ring in the extended Lubin-Tate case. Moreover, based on the expected size of the isogeny classes in the Newton stratum, we conjecture that this lower bound is sharp.
@@ 1.1. The Newton stratum
Throughout this paper, ${\mathbb{F}}_{q}$ will denote the finite field with $q$ elements where $q$ is a power of an odd prime. We will denote the algebraically closed field $\overline{{\mathbb{F}}}_{q}$ by $k$. Let $\mathcal{A}_{g,1}$ denote the moduli space of $g$ dimensional principally polarised abelian varieties in characteristic $p.$ This moduli space is often studied by describing it as a disjoint union of locally closed subvarieties. One way to do this is to introduce a “stratification” using the Newton polygon which is an invariant in characteristic $p$. The Newton stratum $N$ associated to an abelian variety $A,$ consists of the points of $\mathcal{A}_{g,1}$ that have the same Newton polygon as that of $A$. In particular, it contains all the elements in the isogeny class of $A.$ In his paper [Oor04], Frans Oort gives an “almost product” structure to the Newton stratum, which we call a “foliation”. The foliation consists of two “transversal” subvarieties known as the “central leaf” and the “isogeny leaf”. The central leaf contains all the abelian varieties over ${\mathbb{F}}_{q^{n}}$ whose $p$-divisible group is isomorphic to that of $A$ over $k$. In particular, it contains all the abelian varieties that are isogenous to $A$ by prime-to-$p$ isogenies. While the “isogeny leaf” consists of the abelian varieties that are isogenous to $A$ by a certain $p$-power isogeny.
@@
(Isogeny class) Let $A$ be an abelian variety over ${\mathbb{F}}_{q}$. We define $I(A,{\mathbb{F}}_{q^{n}})$, the set of principally polarized abelian varieties $y\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})$ isogenous to $A$ over $k$. We call this the isogeny class of $A$ defined over ${\mathbb{F}}_{q^{n}}.$
1.1.1. Estimating the set $I(A,{\mathbb{F}}_{q^{n}})$
Let $c$ and $I$ denote the central leaf and the isogeny leaf of $A$ respectively. We know that the number of Weil $q^{n}$-numbers with a fixed Newton polygon is roughly $O(q^{n\dim c/2})$. This is computed in [ST18]. As one should expect, the size of $I(A,{\mathbb{F}}_{q^{n}})$ should be $O(q^{n(\dim c/2+\dim I)})$ to account for roughly $q^{n\dim N}$ points of $N({\mathbb{F}}_{q^{n}})$. The dimensions of the leaves can be computed explicitly using the Newton polygon of $A$ as proved by Oort in [Oor09].
1.1.2. Previous cases
The question of counting isogeny classes is trivial for supersingular abelian varieties. Indeed, all the abelian varieties in the supersingular Newton stratum are isogenous over $k$. Hence the Newton stratum is the entire isogeny class in this case. The lower bounds of $I(A,{\mathbb{F}}_{q^{n}})$ for the ordinary and the almost ordinary cases is worked out in [ST18] and [OS20] respectively. In another paper [BF23] with Yu Fu, we estimate isogeny classes of certain abelian varieties with real multiplication. We provide upper and lower bounds in this case. The second author of the paper, has work in this direction as well. For example see [Fu23] for the case of a product of a supersingular elliptic curve with an ordinary one.
In this paper, we further generalise the work for the Newton stratum of extended Lubin-Tate type.
@@ 1.1.2. Previous cases
A $p$-divisible group ${\mathscr{G}}$ is of extended Lubin-Tate type if its local-local part (or its dual) is trivial or dimension $1$
In the case of $p$-divisible groups of principally polarised abelian varieties, the non-trivial case corresponds to when the $p$-divisible group decomposes as a dimension one and a co-dimension one $p$-divisible group. We note that in this case, the Newton polygon of ${\mathscr{G}}$ has slope $\frac{1}{g}$ and $\frac{g-1}{g}.$
1.1.3. The main theorem
Let $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ be a geometrically simple, abelian variety with Weil $q$-number $\alpha$. Moreover, let $A[p^{\infty}]$ be its $p$-divisible group which we assume to be of extended Lubin-Tate type. Furthermore, we assume that $K={\mathbb{Q}}(\alpha)$ is the endomorphism algebra of $A\otimes{\mathbb{F}}_{q^{n}}$ such that $K\otimes{\mathbb{Q}}_{p}$ is unramified. Let $N(A,{\mathbb{F}}_{q^{n}})$ denote the size of the set $I(A,{\mathbb{F}}_{q^{n}})$. In this paper, we prove the lower bound for $N(A,{\mathbb{F}}_{q^{n}})$ given in the following main result of this paper.
@@ 1.1.3. The main theorem
For a positive density set of positive integers $n$, we have the following lower bound
@@
For a positive density set of positive integers $n$, we have the following estimate
$N(A,{\mathbb{F}}_{q^{n}})=q^{n(\frac{(g+1)(g-2)}{4}+1)+o(1))}.$
The dimension of the Newton stratum is $\frac{g(g-1)}{2}$ and that of the central leaf $c(A)$ is $\frac{(g+1)(g-2)}{2}$. Therefore, the bound in Theorem 1.3 agrees with the expected estimate.
As mentioned in [ST18], one way to prove the upper bound would be to estimate the class groups of all the endomorphisms rings of the isogeny class. We might do this by finding a parametrisation of non maximal orders of $K$ containing the smallest possible endomorphism ring. This seems out of reach. However, methods involving orbital integrals have provided upper bounds in certain cases as well. See for example [AC02].
1.1.4. Comparison of methods from previous cases
@@ 1.2.1. Notation
Let $W({\mathbb{F}}_{q})$ denote the Witt vectors over ${\mathbb{F}}_{q}.$ Let $L=W({\mathbb{F}}_{q})[1/p]$ be the unramified extension of ${\mathbb{Q}}_{p}$ with residue field ${\mathbb{F}}_{q}$, while we let $L^{\prime}$ denote $W(k)[1/p]$, the maximal unramified extension of ${\mathbb{Q}}_{p}.$ Let $v$ be the unramified place of $L^{\prime}$ over $p.$
1.2.2. Strategy
@@
As a first step, we note that there are only finitely many abelian varieties over $L^{\prime}$ that reduce to $A$ modulo $v$. Indeed, the number of such abelian varieties over $L^{\prime}$ with endomorphism algebra $K$, will depend on the number of CM-types of $K$. There are only finitely many such. Furthermore, once we fix a CM type on $K$, we get an injective map from the isogeny class over $L^{\prime}$ to the isogeny class in char $p.$ This is proved in section 3, Proposition 3.1. Therefore, we fix an abelian variety $\widetilde{A}$ over $L$ and a CM-type $(K,\Phi)$ on its endomorphism algebra. Our strategy will be to count the number of abelian varieties over $L^{\prime}$ with the endomorphism ring $R$ contained in $K$ characterised by the smallest local conditions. We compute the local conditions in Section 4. We then restrict our count to the set $I(\widetilde{A},L^{\prime}).$ That is, we count the number of principally polarised abelian varieties isogenous to $\widetilde{A}$ with endomorphism ring $R.$ This is done in section 6 and 8. Note that this count is being done in characteristic 00. By the work of Kisin-Lam-Shankar-Srinivasan [KLSS] we know that the number of abelian varieties over ${\mathbb{F}}_{q^{n}}$ having CM lifts lie in finitely many central leaves. See Theorem $1.3$ in [KLSS] for a precise statement. In fact, in this case taking the special fibre of abelian varieties over $W({\mathbb{F}}_{q^{n}})$ will land in exactly one central leaf (see sub-section 5.2). In order to get the final count, we will need to multiply the count in characteristic zero with $q^{\dim I.}$ It is therefore sufficient to prove that $N(\widetilde{A},L^{\prime})\geq q^{\dim c/2}.$
1.2.3. The unpolarised count
@@
In order to make use of the theory of moduli spaces in characteristic $p$, we need to restrict our question to principally polarised abelian varieties. Indeed, in the polarised case, the completion of $\mathcal{A}_{g,1}$ over ${\mathbb{F}}_{p}$ at a point $A_{0}$, denoted by $\mathcal{A}_{g,1}^{\backslash A_{0}}=\operatorname{spf}(R)$ represents the deformation functor of $A_{0}$ as a formal scheme. If we consider the universal deformation $\mathscr{A_{0}}$ over $\operatorname{spf}(R)$, then that can be algebraised to give us an abelian scheme over $\operatorname{Spec}R$ since $\mathscr{A}$ is principally polarised. The (polarised) central leaf of $A_{0}$ is defined to be all $s\in\operatorname{Spec}(R)$ such that the $p$-divisible group $\mathscr{G}_{s}\otimes k$ of the point is isomorphic to that of $A_{0}$ over $k.$ In the unpolarised case, the universal deformation $\mathscr{A}^{\operatorname{np}}$ over $\operatorname{spf}(S)$ of $A_{0}$ need not algebraise. However, its $p$-divisible group over $\operatorname{spf}(S)$ comes from a $p$-divisible group $\mathscr{G}^{\operatorname{np}}$ over $\operatorname{Spec}(S)$. The notion of (unpolarised) central leaf still makes sense. Suppose $x_{0}$ is the closed point of $\operatorname{Spec}(S)$ that represents $A_{0}$. Then the (unpolarised) central leaf of $A$ is the set of all points $s\in S$ such that $\mathscr{G}_{s}^{\operatorname{np}}\otimes k$ is isomorphic to the $p$-divisible group of the special fibre $\mathscr{G}^{\operatorname{np}}_{x_{0}}$ over $k$. This represents a formal object in the local deformation space of $A_{0}$. The dimension of the unpolarised central leaf is also determined by the Newton polygon of $A_{0}$ and can be computed using the methods outlined in [Oor09].
Now, without referring to moduli spaces, we can ask for the count of isogeny classes of abelian varieties over finite fields without any conditions of polarisation. This count in characteristic zero will be the same as computed in subsection 7.0.2 before we take polarisation into account. This is recorded in corollary 7.2. We note that analogous to the case of polarisation, the lower bound for this estimate is in fact of order $\mathcal{O}(q^{\dim c_{\operatorname{np}}/2})$ where $c_{\operatorname{np}}$ is the unpolarised central leaf in the local deformation space of $A.$ Hence, the count over finite fields will be greater than or equal to $q^{n(\dim(c_{\operatorname{np}}/2+I_{\operatorname{np}})+o(1))}$ where $I_{\operatorname{np}}$ is the unpolarised isogeny leaf of the $p$-divisible group of $A_{0}$.
2. Acknowledgements
@@
I am grateful to my advisor, Ananth Shankar for suggesting this project and for many helpful discussions regarding this paper. I also thank Asvin G. for a helpful conversation regarding Lemma 3.2 of the paper. I was partially supported by the NSF grant DMS-2100436.
@@ 3. Injectivity of the special fibre map
From this section onwards, we fix a geometrically simple abelian variety $\widetilde{A}$ over $L$ such that $\operatorname{End}^{0}(A)=K.$ Moreover, we assume that $\widetilde{A}$ has good reduction at $p.$ By choosing an embedding of $L$ into ${\mathbb{C}},$ every such abelian variety in characteristic 00 induces a CM type $\Phi$ on $K$. Since $\Phi$ determines the isogeny class of $\widetilde{A}$ over an algebraically closed field of characteristic 00 (see Proposition 1.5.4.1 in [CCO]), we work with the corresponding CM type on $K.$ As $\widetilde{A}$ is simple, the CM type $\Phi$ is primitive. The purpose of this section is to prove the following proposition.
Proposition 3.1.
Let $\widetilde{A}$ and $\widetilde{B}$ be abelian varieties over $L^{\prime}$ with good reduction and the same CM type $(K,\Phi).$ Let $\mathscr{A}$ and $\mathscr{B}$ denote their Néron models respectively over $W(k).$ We assume that the special fibre of both the schemes is $A.$ Then $\widetilde{A}$ and $\widetilde{B}$ are isomorphic.
The following two results are technical lemmas that will help us build a method to prove the proposition. We put together these results in 3.0.1 to give the overall idea of the proof. The subsequent results will then help us prove the proposition.
@@
Let $\widetilde{A}$ and $\widetilde{B}$ are abelian varieties defined over $L^{\prime}$ such that they have good reduction at a place $v$ of $L^{\prime}$. Suppose that they induce the same CM-type over the algebraically closed field ¯$\overline{\mathbb{Q}}_{p}$. That is, they are isogenous over ¯$\overline{\mathbb{Q}}_{p}$. Then they are isogenous over $L^{\prime}$.
@@
Let $\widetilde{A}$ and $\widetilde{B}$ be isogenous over a finite extension $F^{\prime}$ of $L^{\prime}$. Since $\widetilde{A}$ and $\widetilde{B}$ both have good reduction, we see that $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})\hookrightarrow\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$. Let $I_{\overline{v}}$ denote the inertia group of $v$ where $\overline{v}$ is the extension of the place $v$ to the separable closure of $L^{\prime}$. We note that $I_{\overline{v}}$ acts trivially on $\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$ and hence on $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})$ as well. Therefore, $\widetilde{A}$ and $\widetilde{B}$ are in fact isogenous over $L^{\prime}.$
@@
Suppose that $\widetilde{A}$ is an abelian variety over $L^{\prime}$ with good reduction. Let $A$ denote its special fibre. Then $\operatorname{End}\widetilde{A}=\operatorname{End}A.$

### v2

# size of isogeny classes of abelian varieties of lubin-tate type

@@ Abstract.
We prove a lower bound for the size of the isogeny class of a simple abelian variety over a finite field with commutative endomorphism ring in the Lubin-Tate case. Moreover, based on the expected size of the isogeny classes in the Newton stratum, we conjecture that this lower bound is sharp.
Key words and phrases:
Abelian varieties, finite fields, complex multiplication, isogeny class, Newton stratum
2000 Mathematics Subject Classification:
11G15 (Primary), 11G10 (Secondary).
@@ 1.1. The Newton stratum
Throughout this paper, we denote ${\mathbb{F}}_{q}$ to be the finite field with $q$ elements where $q$ is a power of an odd prime. For any positive integer $n$, we denote by ${\mathbb{F}}_{q^{n}}$ to be the unique finite field extension of ${\mathbb{F}}_{q}$ of degree $n$. We will denote the algebraically closed field $\overline{{\mathbb{F}}}_{q}$ by $k$. Let $\mathcal{A}_{g,1}$ denote the moduli space of $g$ dimensional principally polarised abelian varieties in characteristic $p.$ This moduli space is often studied by describing it as a disjoint union of locally closed subvarieties. One way to do this is to introduce a “stratification” using the Newton polygon which is an invariant in characteristic $p$. Let $A$ be a principally polarised abelian variety over ${\mathbb{F}}_{q}$ and let $N(A)$ be the associated Newton stratum. We know that this is locally closed subvariety of $\mathcal{A}_{g,1}$. Then $N(A)({\mathbb{F}}_{q^{n}})$ consists of the points of $\mathcal{A}_{g,1}$ defined over ${\mathbb{F}}_{q^{n}}$ that have the same Newton polygon as that of $A$. In particular, it contains all the elements in the isogeny class of $A$ as defined below.
@@
(Isogeny class) We define $I(A,{\mathbb{F}}_{q^{n}})$, to be the set of isomorphism classes of principally polarized abelian varieties $y\in\mathcal{A}_{g,1}({\mathbb{F}}_{q^{n}})$ isogenous to $A$ over $k$. We call this the isogeny class of $A$ defined over ${\mathbb{F}}_{q^{n}}.$
In his paper [Oor04], Frans Oort gives an “almost product” structure to the Newton stratum, which we call a “foliation”. The foliation consists of two “transversal” subvarieties known as the “central leaf” and the “isogeny leaf”. The central leaf contains all the abelian varieties over ${\mathbb{F}}_{q^{n}}$ whose $p$-divisible group is isomorphic to that of $A$ over $k$. In particular, it contains all the abelian varieties that are isogenous to $A$ by prime-to-$p$ isogenies. While the “isogeny leaf” consists of the abelian varieties that are isogenous to $A$ by a certain $p$-power isogeny. Using this structure, Shankar and Tsimerman (See [ST18], Section 2) find heuristics for the size of isogeny classes of abelian varieties. We briefly describe this.
1.2. Heuristics for the size of the isogeny class.
Let $c(A)$ and $\iota(A)$ denote the central leaf and the isogeny leaf of $A$ respectively.
@@
(Weil $q$-number of an abelian variety) Let $A$ be as above and $\alpha_{A}$ be its geometric Frobenius. Further assume that $A$ is simple. Then we know that $\alpha_{A}$ is an algebraic integer such that for every embedding $\psi:{\mathbb{Q}}(\alpha_{A})\hookrightarrow{\mathbb{C}}$, we have that $|\psi(\alpha_{A})|=\sqrt{q}.$ Such an algebraic integer is called a Weil $q$-number. By Honda Tate theory we know that the isogeny classes of simple abelian varieties are in bijection with Weil $q$-numbers under $A\mapsto\alpha_{A}$.
We know that the number of Weil $q^{n}$-numbers with a fixed Newton polygon is roughly $O(q^{n\dim c(A)/2})$ which gives us an estimate for the number of ${\mathbb{F}}_{q^{n}}$ isogeny classes in the Newton stratum. This is computed in [ST18], Lemma 2.2. As one should expect, the size of $I(A,{\mathbb{F}}_{q^{n}})$ should be $O(q^{n(\dim c(A)/2+\dim\iota(A))})$ to account for roughly $q^{n\dim N}$ points of $N({\mathbb{F}}_{q^{n}})$. The dimensions of the central leaves and Newton strata can be computed explicitly and only depends on the Newton polygon of $A$. This was also proved by a follow up paper by Oort (See [Oor09], Section 5 and 6).
1.3. Previous work
The question of counting isogeny classes is trivial for supersingular abelian varieties. Indeed, all the abelian varieties in the supersingular Newton stratum are isogenous over $k$. Hence the Newton stratum is the entire isogeny class in this case. One of the first results goes back to the upper bound for elliptic curves by Lenstra [Len87] and subsequent work provides mainly lower bounds in higher dimensions. The lower bounds of $I(A,{\mathbb{F}}_{q^{n}})$ for the ordinary and the almost ordinary cases is worked out in [ST18] and [OS20] respectively. In another paper [BF23] with Yu Fu, we estimate isogeny classes of certain abelian varieties with real multiplication. We provide upper and lower bounds in this case. The second author of the paper, has work in this direction as well. For example, see [Fu23] for the case of a product of a supersingular elliptic curve with an ordinary one.
In this paper, we further generalise the work for the Newton stratum of Lubin-Tate type.
Definition 1.3.
A $p$-divisible group is of Lubin-Tate type if the dimension of its local-local part or that of its dual is atmost one.
In the case of $p$-divisible groups of principally polarised abelian varieties, the non-trivial case corresponds to when the $p$-divisible group decomposes as a dimension one and a co-dimension one $p$-divisible group. We note that in this case, its Newton polygon has slope $1/g$ and $(g-1)/g.$
1.4. The main theorem
Let $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ be a geometrically simple, abelian variety with Weil $q$-number $\alpha_{A}$. For the sake of convenience, we drop of the subscript and the dependence on $A$ will be implicit. Let the $p$-divisible group $A[p^{\infty}]$ of $A$ be of Lubin-Tate type. Let $K={\mathbb{Q}}(\alpha)$ be the the endomorphism algebra of $A\otimes k$. Due to the condition on the Newton polygon, the tensor product $K\otimes{\mathbb{Q}}_{p}$ is a direct sum of two field extensions of ${\mathbb{Q}}_{p}$ of degree $g$ which we assume to be unramified.
For any positive integer $n$, let $N(A,{\mathbb{F}}_{q^{n}})$ denote the size of the set $I(A,{\mathbb{F}}_{q^{n}})$. We prove the lower bound for $N(A,{\mathbb{F}}_{q^{n}})$ given in the following main result of this paper.
@@ 1.4. The main theorem
For a set of positive integers $n$ with positive natural density, we have the following lower bound
@@
For a set of positive integers $n$ with positive natural density, we have the following estimate
$N(A,{\mathbb{F}}_{q^{n}})=q^{n(\frac{(g+1)(g-2)}{4}+1+o(1))}.$
The dimension of the Newton stratum is $g(g-1)/2$ and that of the central leaf $c(A)$ is $(g+1)(g-2)/2$. Therefore, the bound in Theorem 1.4 agrees with the expected estimate.
As mentioned in [ST18], one way to prove the upper bound would be to estimate the class groups of all the endomorphisms rings of the isogeny class. We might do this by finding a parametrisation of non maximal orders of $K$ containing the smallest possible endomorphism ring. This seems out of reach. However, methods involving orbital integrals have provided upper bounds in certain cases as well. See for instance [AC02].
1.5. Comparison of methods from previous cases
@@ 1.6. Overall strategy
Let $W(-)$ denote the functor of Witt vectors. Let $L=W({\mathbb{F}}_{q})(1/p)$ be the unramified extension of ${\mathbb{Q}}_{p}$ with residue field ${\mathbb{F}}_{q}$, and $L^{\prime}$ be $W(k)(1/p)$, the maximal unramified extension of ${\mathbb{Q}}_{p}.$ Let $v$ be the unramified place of $L^{\prime}$ over $p.$
Counting in characteristic zero
We count in characteristic zero (over $L^{\prime})$ as follows. We fix $A\in\mathcal{A}_{g,1}({\mathbb{F}}_{q})$ as in Theorem 1.4 and an abelian variety $\widetilde{A}$ over $L$ that reduces to $A$ mod $p$. In addition, we fix a CM-type $(K,\Phi)$ on the endomorphism algebra of $\widetilde{A}$.
As a first step, we note that there are only finitely many abelian varieties over $L^{\prime}$ that reduce to $A$ (base changed to $L^{\prime}$) modulo $v$. Indeed, the number of such abelian varieties over $L^{\prime}$ with endomorphism algebra $K$, will depend on the number of CM-types of $K$. There are only finitely many such. Furthermore, once we fix a CM type on $K$, we get an injective map from the isogeny class over $L^{\prime}$ to the isogeny class in char $p.$ This is proved in Section 2, Proposition 2.7. Therefore, it is enough to count the number of abelian varieties over $L^{\prime}$ with a fixed CM type. Let $n$ be any positive integer. Now to get a lower bound on $I(A,{\mathbb{F}}_{q^{n}})$, we reduce the problem to counting the set $I(A,L^{\prime})$ whose endomorphism algebra $K$ contains a certain fixed endomorphism ring $R_{n}$. We then give an injection from $Cl(R_{n})$, the class group of $R_{n}$ to $I(A,L^{\prime})$. We note that the ring $R_{n}$ depends on $n$ and is characterized by certain local conditions at the prime $p$, that we compute in Section 3. It is crucial to take into account these conditions as they help in estimating the class group of $R_{n}$. Moreover, these local conditions ensure that all the abelian varieties in the set $I(A,L^{\prime})$ with endomorphism ring $R_{n}$ descend to $W({\mathbb{F}}_{q^{n}})(1/p)$ (See Lemma 4.1), consequently giving us a sharp lower bound for $I(A,{\mathbb{F}}_{q^{n}})$ once we reduce mod $p.$
Reducing mod $p$
@@
By the work of Kisin-Lam-Shankar-Srinivasan we know that the number of abelian varieties over ${\mathbb{F}}_{q^{n}}$ having CM lifts lie in finitely many central leaves. See Theorem $1.3$ in [KLSS25] for a precise statement. In fact, in this case taking the special fibre of abelian varieties over $W({\mathbb{F}}_{q^{n}})$ will land in exactly one central leaf (See Section 4.2). It is therefore sufficient to prove that $N(\widetilde{A},L^{\prime})\geq q^{\dim c(A)/2}.$ In order to get the final count, we will need to multiply the count in characteristic zero with $q^{\dim\iota(A)}$ thus proving the heuristic in Section 1.2.
1.7. Organisation of the paper
In Section 2 we prove that the reduction mod $p$ map from the isogeny class in characteristic zero to finite fields is injective. In Section 3 and 4, we compute the local conditions for the endomorphism ring that we use to count and certain properties of abelian varieties with that particular endomorphism ring. In Section 6, we give an injection between the isogeny class and the class of the order. In Section 5, we find bounds for the (unpolaised) isogeny class by estimating the class group of the order. Finally, Section 7 is dedicated to taking polarisation into account for the count.
1.8. The unpolarised count
@@
In order to make use of the theory of moduli spaces in characteristic $p$, we need to restrict our question to principally polarised abelian varieties. Indeed, in the polarised case, the completion of $\mathcal{A}_{g,1}$ over ${\mathbb{F}}_{p}$ at an ${\mathbb{F}}_{q}$-point $A$, denoted by $\mathcal{A}_{g,1}^{\backslash A}=\operatorname{spf}(Y)$ represents the deformation functor of $A$ as a formal scheme. If we consider the universal deformation $\mathscr{A}$ over $\operatorname{spf}(Y)$, then that can be algebraised to give us an abelian scheme over $\operatorname{Spec}(Y)$ since $\mathscr{A}$ is principally polarised. The polarised central leaf of $A$ is defined to be all $s\in\operatorname{Spec}(Y)$ such that the $p$-divisible group $\mathscr{G}_{s}\otimes k$ of the point is isomorphic to that of $A$ over $k.$ In the unpolarised case, the universal deformation $\mathscr{A}^{\operatorname{np}}$ over $\operatorname{spf}(X)$ of $A$ need not algebraise. However, its $p$-divisible group over $\operatorname{spf}(X)$ comes from a $p$-divisible group $\mathscr{G}^{\operatorname{np}}$ over $\operatorname{Spec}(X)$. The notion of unpolarised central leaf still makes sense. Suppose $x_{0}$ is the closed point of $\operatorname{Spec}(X)$ that represents $A$. Then the unpolarised central leaf of $A$ is the set of all points $s\in S$ such that $\mathscr{G}_{s}^{\operatorname{np}}\otimes k$ is isomorphic to the $p$-divisible group of the special fibre $\mathscr{G}^{\operatorname{np}}_{x_{0}}$ over $k$. This represents a formal object in the local deformation space of $A$. The dimension of the unpolarised central leaf is also determined by the Newton polygon of $A$ and can be computed using the methods outlined in [Oor09], Section 4. Now, without referring to moduli spaces, we can ask for the count of isogeny classes of abelian varieties over finite fields without any conditions of polarisation. This count in characteristic zero will be the same as computed in subsection 6.2 before we take polarisation into account. This is recorded in corollary 6.2. We note that analogous to the case of polarisation, the lower bound for this estimate is in fact of order $\mathcal{O}(q^{\dim c_{\operatorname{np}}/2})$ where $c_{\operatorname{np}}$ is the unpolarised central leaf in the local deformation space of $A.$ Hence, the count over finite fields will be greater than or equal to $q^{n(\dim(c_{\operatorname{np}}/2+\iota_{\operatorname{np}})+o(1))}$ where $\iota_{\operatorname{np}}$ is the unpolarised isogeny leaf of the $p$-divisible group of $A$.
Acknowledgements
@@
I am grateful to Ananth Shankar for suggesting this project and for many helpful discussions regarding this paper. I also thank Asvin G. for a helpful conversation regarding Lemma 2.1 of the paper. I was partially supported by the NSF grant DMS-2100436.
@@ 2. Injectivity of the special fibre map
From this section on wards, we fix $A\in\mathcal{A}_{g,1}$ over ${\mathbb{F}}_{q}$ as in the main Theorem 1.4. Let $\widetilde{A}$ abelian variety $L=W({\mathbb{F}}_{q})(1/p)$ be reducing to $A$ over ${\mathbb{F}}_{q}$. We assume both $A$ and $\widetilde{A}$ to be geometrically simple. Let $\operatorname{End}^{0}(A)=K=\operatorname{End}^{0}(\widetilde{A})$. Choosing an embedding of $L$ into ${\mathbb{C}},$ induces a CM type $\Phi$ on $K$. We fix one such CM type $\Phi$ on $K$ that determines the isogeny class of $\widetilde{A}$ over an algebraically closed field of characteristic 00 (See Proposition 1.5.4.1 in [CCO14]). In fact we have the following lemma:
@@
Let $\widetilde{A}$ (base changed to $L^{\prime}=W(k)(1/p))$ be as above and $\widetilde{B}$ be another abelian variety defined over $L^{\prime}$ such that they have good reduction at a place $v$ of $L^{\prime}$. Suppose that they induce the same CM-type over the algebraically closed field ¯$\overline{\mathbb{Q}}_{p}$. That is, they are isogenous over ¯$\overline{\mathbb{Q}}_{p}$. Then they are isogenous over $L^{\prime}$.
@@
Let $\widetilde{A}$ and $\widetilde{B}$ be isogenous over a finite extension $F^{\prime}$ of $L^{\prime}$. Since $\widetilde{A}$ and $\widetilde{B}$ both have good reduction, we see that $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})\hookrightarrow\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$. Here $\widetilde{A}_{-}$ denotes the abelian variety base changed to the particular field. Let $I_{\overline{v}}$ denote the inertia group of $v$ where $\overline{v}$ is the extension of the place $v$ to the separable closure of $L^{\prime}$. We note that $I_{\overline{v}}$ acts trivially on $\operatorname{Hom}(\widetilde{A}_{k},\widetilde{B}_{k})$ and hence on $\operatorname{Hom}(\widetilde{A}_{F^{\prime}},\widetilde{B}_{F^{\prime}})$ as well. Therefore, $\widetilde{A}$ and $\widetilde{B}$ are isogenous over $L^{\prime}.$
We further observe that the reduction map preserves the endomorphism orders.
@@
Let $\widetilde{A}$ and $A$ be as above. Then $\operatorname{End}\widetilde{A}=\operatorname{End}A.$